OpenStudy (anonymous):

determine the type number solution 3x^2+6x+4=0 a.1 real solution b.one real and 1imaginary solution c. 2 imaginary solution d. 2 real solution

OpenStudy (anonymous):

Sqrt[b^2 - 4ac]

OpenStudy (anonymous):

that is 6^2-4*3*4=36-48=-12 sqrt of -12 is an imaginary number, so there are 2 imaginary solutions

OpenStudy (anonymous):

by the way of there is an equation like x^n+x^(n-1)+...+1=0 it will always have n number of solutions if we allow complex numbers